OpenStudy (anonymous):

The x intercept of a line is-5 and the y intercept is -2 what is the equation of the line?

Parth (parthkohli):

That's another way to say that the line passes through the points (-5,0) and (0,-2)

OpenStudy (anonymous):

After that, you find the slope of the line. \[y = mx + b\] where \(m\) is the slope of the line and \(b\) is the y-intercept.

OpenStudy (anonymous):

\[\large \frac{x}{a}+\frac{y}{b}=1\] a = x-intercept b = y-intercept. This is the intercept-intercept form of the equation of a line.

OpenStudy (anonymous):

the equation of the line is Y=2/5x+b..... Plot the points given by the x intecept (5,0) and the y intercept (0,-2) on a graph or insert the numbers in the slope formula m=y2-y1/x2-x1 either way you get a slope of 2/5 and you put that in the Y intercept form.

OpenStudy (anonymous):

Thank you!!